Please help me evaluate this expression without using a calculator:
(-32)^3/5

Answer by lwsshak3(6522)

Please help me evaluate this expression without using a calculator:
(-32)^3/5
..
(-32)^(3/5)
To solve this expression, take the 5th root of -32, then cube the result.
5th root of -32=-2
-2^3=-8 (ans)
Note:For fractional exponents, the denominator is the root and the numerator is the power.Usually, it is best to take the root first then raise it to the power, although it can be the other way around, power first then root second.